Homework Assignment:  

Determine the present value of each perpetuity 

Perpetuity 

Annual Amount 

Discount rate 

$20,000 

8% 

100,000 

10 

3,000 

60,000 




1
Expert's answer
2021-12-29T12:30:24-0500

Formula to calculate present value of perpetuity is:

"PV=\frac{CF}{r}"

Where;

PV = present value

CF = annual cash flow

r = discount rate


"PV(A)=\frac{20,000}{0.08}=\$250,000"


"PV(B)=\frac{100,000}{0.10}=\$1,000,000"


"PV(C)=\frac{3,000}{0.06}=\$50,000"


"PV(D)=\frac{60,000}{0.05}=\$1,200,000"